OpenStudy (stupidzoe):

How do I solve this ?? (X^3+3x^2-6x+2)/(x-1)

OpenStudy (deadpooi):

x3+3x2−6x+2 x−1 = x3+3x2−6x+2 x−1 = (x−1)(x2+4x−2) x−1 =x2+4x−2